OpenStudy (anonymous):

Hannah has a jar filled with 150 coins made up of quarters and nickels. The total value of the coins is $28.30. Let x represent the number of quarters in the jar. Which equation can be used to find the number of quarters?

OpenStudy (anonymous):

.05n = (28.30 - .25x) Multiply eq by 20 n = (566 - 5x) : find x using this equation 150 - x = 566 - 5x -x + 5x = 566 - 150 4x = 416 x = 104 quarters